OpenStudy (anonymous):

Explain how you could use the Pythagorean Theorem to find the distance between the points (-4,-3) and (1,1) . Find the distance. You can type"^2" to indicate "squared" and "sqrt" to indicate "square root." (Examples:5^2=5^2 and sqrt (25)=√25

OpenStudy (binary3i):

 AC=(1-(-4))=5 BC=(1-(-3))=4 now you can use pythagorean